OpenStudy (anonymous):

(Algebra II) x/a - 5 = b Solve for X, find any restrictions. Please help!!

jimthompson5910 (jim_thompson5910):

Is the equation \[\Large \frac{x}{a} - 5 = b\] OR \[\Large \frac{x}{a - 5} = b\] ??

OpenStudy (anonymous):

It's the first equation :)

OpenStudy (anonymous):

@jim_thompson5910

jimthompson5910 (jim_thompson5910):

start by adding 5 to both sides \[\Large \frac{x}{a} - 5 = b\] \[\Large \frac{x}{a} -5+5= b+5\] \[\Large \frac{x}{a} = b+5\]

jimthompson5910 (jim_thompson5910):

what's next?

OpenStudy (anonymous):

@jim_thompson5910 I wanna say multiply each side by something, I think multiply by x or a on both sides but I'm not sure.

jimthompson5910 (jim_thompson5910):

you want to keep 'x' on the left side

jimthompson5910 (jim_thompson5910):

so you have to multiply both sides by 'a'

OpenStudy (anonymous):

@jim_thompson5910 Ahhh okay :)!

jimthompson5910 (jim_thompson5910):

\[\Large \frac{x}{a} = b+5\] \[\Large a\left(\frac{x}{a}\right) = a(b+5)\] \[\Large \cancel{a}\left(\frac{x}{\cancel{a}}\right) = a(b+5)\] \[\Large x = a(b+5)\]

jimthompson5910 (jim_thompson5910):

Optionally, you can distribute the right side, but it's not necessary

OpenStudy (anonymous):

@jim_thompson5910 I see it now, thank you very much!! :) Helped a lot. I'll try to apply that to my other problems as well! :)

jimthompson5910 (jim_thompson5910):

alright great, yw